# Offline Safe Reinforcement Learning Using Trajectory Classification

### v4

@@ 5.3 Results
In Figure 4, we present the results of varying $x\%$ while keeping $y\%$ fixed in two tasks. On the x-axis, the notation $+x\%-y\%$ represents the experimental setting. In AntCircle task, the normalized reward remains consistent across difference $x\%$ selections, while the normalized cost decreases as $x\%$ increases. In MetaDrive-mediummean task, remains relatively unchanged with varying $x\%$. The highest normalized reward is observed when $x\%$ is set to $75\%$, although the differences are minor. Additionally, we conducted experiments with varying $y\%$ while keeping $x\%$ fixed, as shown in Figure 5. Similarly, altering $y\%$ results in only minor changes in both test tasks. These findings suggest that TraC is robust and exhibits little sensitivity to different selections of $x\%$ and $y\%$.

### v5

@@ 5.3 Results
In Figure 4, we present the results of varying $x\%$ while keeping $y\%$ fixed in two tasks. On the x-axis, the notation $+x\%-y\%$ represents the experimental setting. In AntCircle task, the normalized reward remains consistent across difference $x\%$ selections, while the normalized cost decreases as $x\%$ increases. We also conducted experiments with varying $y\%$ while keeping $x\%$ fixed, as shown in Figure 5. Altering $y\%$ results in minor changes in normalized reward, with the lowest normalized cost obtained when $y\%$ is set to $0\%$. While in MetaDrive-mediummean task, both the normalized reward and cost decrease as either $x\%$ or $y\%$ increases. These findings suggest that TraC is generally robust to varying selections of $x\%$ and $y\%$ though further fine-tuning for specific tasks may yield improved performance.